# Generalizable Domain Adaptation for Sim-and-Real Policy Co-Training

### v2

@@ 4.2 Unbalanced Optimal Transport for Robust Alignment
The standard OT formulation (Eq. 1) enforces strict marginal constraints, requiring all mass from the source distribution to be transported to the target and vice-versa. This is problematic in our sim-to-real setting primarily due to significant data imbalance ($|D_{src}|\gg|D_{tgt}|$) and the partial overlap between the state spaces covered by $D_{src}$ and $D_{tgt}$ (Section 3.1). Standard OT would either distort the latent space by forcing many-to-few mappings or create spurious alignments between non-corresponding states. To address these challenges, we employ Unbalanced Optimal Transport (UOT) chizat2018scaling ; fatras2021unbalanced . UOT relaxes the hard marginal constraints of OT by introducing regularization terms that penalize deviations, thereby allowing for partial mass transport. This enables UOT to selectively align subsets of the distributions that are most similar according to the ground cost, while effectively down-weighting or ignoring the transport for dissimilar or unmatched portions.
UOT Loss Formulation. Consider a mini-batch of $N_{batch}$ source samples $\{(o^{i}_{src},x^{i}_{src})\}_{i=1}^{N_{batch}}$ and $N_{batch}$ target samples $\{(o^{j}_{tgt},x^{j}_{tgt})\}_{j=1}^{N_{batch}}$. Let their empirical distributions in the joint $(f_{\phi}(o),x)$ space be $\hat{\mu}_{src}$ and $\hat{\mu}_{tgt}$. Our UOT loss, $L_{UOT}(f_{\phi})$, is based on the Kantorovich formulation with entropic regularization and KL-divergence penalties for marginal relaxation:
@@
Here, $\hat{C}_{\phi}$ is the $N_{batch}\times N_{batch}$ ground cost matrix, where each element $(\hat{C}_{\phi})_{ij}$ is computed using the joint ground cost described in Sec. 4.1. The term $\Pi$ is the transport plan; $\epsilon>0$ is the entropic regularization strength with $\Omega(\Pi)=\sum_{i,j}\Pi_{ij}\log\Pi_{ij}$ being the entropy, facilitating efficient solution via algorithms like Sinkhorn-Knopp cuturi2013sinkhorn ; $\tau>0$ controls the penalty for deviating from the batch marginals $\mathbf{p}$ and $\mathbf{q}$ (typically uniform); and $\text{KL}(\cdot||\cdot)$ denotes the Kullback-Leibler divergence.
@@ 4.3 Temporally Aligned Sampling for Effective Mini-Batch Learning
To address this, we introduce a temporally aligned sampling strategy designed to construct mini-batches with a higher density of meaningfully corresponding state-pairs. Our strategy leverages trajectory-level similarity as a heuristic. We first quantify similarity between source trajectories $\{\xi^{k}_{src}\}\subset D_{src}$ and target trajectories $\{\xi^{l}_{tgt}\}\subset D_{tgt}$ using Dynamic Time Warping (DTW) gold2018dynamic on their respective proprioceptive state sequences $\{x_{t}\}$. The resulting normalized DTW distance, $\bar{d}(\xi^{k}_{src},\xi^{l}_{tgt})=d_{DTW}(\xi^{k}_{src},\xi^{l}_{tgt})/\max(|\xi^{k}_{src}|,|\xi^{l}_{tgt}|)$, reflects overall behavioral similarity. To turn these distances into sampling weights, we apply a softplus-based transformation: $w(\xi^{k}_{src},\xi^{l}_{tgt})=1/(1+e^{10\cdot(\bar{d}(\xi^{k}_{src},\xi^{l}_{tgt})-0.01)})$. Mini-batch construction then proceeds by 1. sampling a pair of trajectories $(\xi_{src},\xi_{tgt})$ with probability biased towards pairs exhibiting high similarity (i.e., low DTW distance) and 2. subsequently sampling individual transition tuples $(o_{src},x_{src},a_{src})$ and $(o_{tgt},x_{tgt},a_{tgt})$ from this selected, behaviorally similar trajectory pair. Fine-grained temporal alignment, such as sampling around DTW-matched time steps, can optionally be employed here. We describe how the UOT loss (Eq. 3) is adapted with this new sampling procedure in the appendix.
@@ 4.4 Joint Co-Training Framework
Putting all components together, our final approach is a joint co-training framework where the visual feature encoder $f_{\phi}$ and the policy $\pi_{\theta}(a|z,x)$ are optimized concurrently. The Unbalanced Optimal Transport loss ($L_{\text{UOT}}$) serves as a regularization term, guiding $f_{\phi}$ to learn domain-invariant and action-relevant latent representations $z=f_{\phi}(o)$, while standard Behavior Cloning (BC) losses drive the policy learning. The overall training objective $L(f_{\phi},\pi_{\theta})=L_{BC}(f_{\phi},\pi_{\theta})+\lambda\cdot L_{\text{UOT}}(f_{\phi})$ combines these components, where $L_{BC}$ represents the combined behavior cloning losses calculated over both source ($D_{src}$) and target ($D_{tgt}$) datasets using a standard imitation loss (e.g., MSE). The hyper-parameter $\lambda>0$ balances feature alignment with policy imitation. The $L_{\text{UOT}}(f_{\phi})$ term is computed as defined in Equation 3, with mini-batches sampled with strategy described in Sec. 4.3. The overall training process is detailed in the appendix.
@@ 5.1 Experiment Setups
Environment setup. For the real-world experiments, we deploy our method on a Franka Emika Panda robotic arm, controlled at 20Hz using joint impedance control. Visual and point cloud observations are captured using an Intel RealSense D435 depth camera. Our simulation environments use the Robosuite zhu2020robosuite simulation framework. We calibrate the camera pose and camera intrinsics in simulation to match those obtained from the real-world setup to reduce the domain gap.
Simulation data. For simulation experiments, we begin by collecting 10 human demonstrations per task. Using MimicGen mandlekar2023mimicgen , we synthesize 200-1000 trajectories in the source domain, covering the full range of initial states (denoted as Source). In the target domain, we divide the reset region into two subregions: one is populated with 10 trajectories for training (denoted as Target), while the other remains completely held out from training (denoted as Target-OOD). This held-out subregion is used to evaluate each method’s generalization under Out-Of-Distribution (OOD) conditions.
@@
Observation modality and domain gaps. We evaluate our approach using two observation modalities: point clouds and RGB images. For point cloud observations, our method and the baselines adapt 3D Diffusion Policy (Ze2024DP3, ) with a PointNet encoder (qi2017pointnet, ). For RGB image observations, we use Diffusion Policy (chi2023diffusionpolicy, ) with a ResNet-18 encoder (he2016deep, ). To evaluate the generalization capabilities of different methods under visual domain shifts in simulation, we introduce several target domain variations: Viewpoint1-Point, Viewpoint3-Point, Perturbation-Point, Viewpoint-Image, and Texture-Image. Descriptions of each domain shift are provided in the appendix.
Baselines. We compare our method against the following baselines: MMD—minimizes the distance between the mean embeddings of source and target data (tzeng2014deep, ); Co-training—trains the model using a mixed batch of source and target domain data, following the strategy proposed by (wei2025empirical, ; maddukuri2025sim, ); Source-only—trains the model exclusively on data from the source domain, which in sim-to-real experiments corresponds to using only simulation data; Target-only—trains the model exclusively on data from the target domain, which in sim-to-real experiments trains with only real world data.
@@ 5.2 Cross-Domain Generalization Results
Our method effectively learns complex manipulation tasks in both simulation and the real world (H1). Experimental results show that our approach consistently matches or outperforms in terms of success rates all baselines across source and target domains in both simulated and real-world settings. On real-world tasks, our method achieves average success rates of 0.73 and 0.77 for image-based and point cloud-based policies, respectively. The Target-only baseline performs well in distribution but fails to generalize under domain shifts. The MMD baseline tzeng2014deep offers limited improvement by aligning global feature statistics, but its coarse alignment often disrupts task-relevant structure and harms source-domain performance. In contrast, by using Unbalanced Optimal Transport, our method performs selective, structure-aware alignment, avoiding spurious matches.
Our method generalizes to target domains only seen in simulation (H2). In Target-OOD scenarios, our method outperforms all baselines, underscoring the value of learning domain-invariant representations (see Tab. 2 and Tab. 4). While Co-training baselines wei2025empirical ; maddukuri2025sim perform well when the target-domain training data overlaps with the evaluation region, they struggle to generalize when this overlap is absent. This limitation is especially evident under large domain shifts. For example, in the real-world BoxInBin and Stack tasks with novel textures or reset poses, our method achieves success rates of 0.7 and 0.4, respectively, using image-based observations—compared to just 0.3 and 0.0 for the Co-training baseline. These results highlight the shortcomings of relying purely on supervised target-domain data without explicitly addressing domain shift.
@@
Our method learns shared latent space that aligns simulation and real data. To better understand how the learned embeddings contribute to generalization, we visualize them using t-SNE van2008visualizing , as shown in Fig. 4(a). Blue and red correspond to features extracted from source and target domain observations, respectively. The left plot shows embeddings from the encoder trained with the Co-training baseline, while the right plot shows embeddings from our method. The visualization reveals that our approach leads to significantly better alignment between source and target distributions, highlighting its ability to learn domain-invariant representations that facilitate robust generalization. We also visualize the transport plan on a randomly sampled batch, along with the corresponding image observations from the source and target domains. As shown in Fig. 10 in the Appendix, the transport plan effectively aligns data points with similar states across domains.
@@ 6 Limitations and Conclusions
The main limitation of our work is that we only address sim-to-real visual observation gaps. Addressing action dynamics gaps remains future work. Because we use MimicGen mandlekar2023mimicgen for automated simulation demonstration generation, we inherit its limitations, namely, our method primarily applies to tasks with prehensile and quasi-static interactions. We require a small number of real-world on-task demonstrations that are aligned with the simulated demonstrations. Future work involves relaxing this requirement, for example, by instead consuming unstructured real-world data, such as play data.

### v3

@@ 4.2 Unbalanced Optimal Transport for Robust Alignment
The standard OT formulation (Eq. 1) enforces strict marginal constraints, requiring all mass from the source distribution to be transported to the target and vice-versa. This is problematic in our sim-to-real setting primarily due to significant data imbalance ($|D_{src}|\gg|D_{tgt}|$) and the partial overlap between the state spaces covered by $D_{src}$ and $D_{tgt}$ (Section 3.1). Standard OT would either distort the latent space by forcing many-to-few mappings or create spurious alignments between non-corresponding states. To address these challenges, we employ Unbalanced Optimal Transport (UOT) Chizat et al. (2018); Fatras et al. (2021). UOT relaxes the hard marginal constraints of OT by introducing regularization terms that penalize deviations, thereby allowing for partial mass transport. This enables UOT to selectively align subsets of the distributions that are most similar according to the ground cost, while effectively down-weighting or ignoring the transport for dissimilar or unmatched portions.
UOT Loss Formulation. Consider a mini-batch of $N_{batch}$ source samples $\{(o^{i}_{src},x^{i}_{src})\}_{i=1}^{N_{batch}}$ and $N_{batch}$ target samples $\{(o^{j}_{tgt},x^{j}_{tgt})\}_{j=1}^{N_{batch}}$. Let their empirical distributions in the joint $(f_{\phi}(o),x)$ space be $\hat{\mu}_{src}$ and $\hat{\mu}_{tgt}$. Our UOT loss, $L_{\text{UOT}}(f_{\phi})$, is based on the Kantorovich formulation with entropic regularization and KL-divergence penalties for marginal relaxation:
@@
Here, $\hat{C}_{\phi}$ is the $N_{batch}\times N_{batch}$ ground cost matrix, where each element $(\hat{C}_{\phi})_{ij}$ is computed using the joint ground cost described in Sec. 4.1. The term $\Pi$ is the transport plan; $\epsilon>0$ is the entropic regularization strength with $\Omega(\Pi)=\sum_{i,j}\Pi_{ij}\log\Pi_{ij}$ being the entropy, facilitating efficient solution via algorithms like Sinkhorn-Knopp Cuturi (2013); $\tau>0$ controls the penalty for deviating from the batch marginals $\mathbf{p}$ and $\mathbf{q}$ (typically uniform); and $\text{KL}(\cdot||\cdot)$ denotes the Kullback-Leibler divergence.
@@ 4.3 Temporally Aligned Sampling for Effective Mini-Batch Learning
To address this, we introduce a temporally aligned sampling strategy designed to construct mini-batches with a higher density of meaningfully corresponding state-pairs. Our strategy leverages trajectory-level similarity as a heuristic. We first quantify similarity between source trajectories $\{\xi^{k}_{src}\}\subset D_{src}$ and target trajectories $\{\xi^{l}_{tgt}\}\subset D_{tgt}$ using Dynamic Time Warping (DTW) Gold and Sharir (2018) on their respective proprioceptive state sequences $\{x_{t}\}$. The resulting normalized DTW distance, $\bar{d}(\xi^{k}_{src},\xi^{l}_{tgt})=d_{\text{DTW}}(\xi^{k}_{src},\xi^{l}_{tgt})/\max(|\xi^{k}_{src}|,|\xi^{l}_{tgt}|)$, reflects overall behavioral similarity. To turn these distances into sampling weights, we apply a softplus-based transformation: $w(\xi^{k}_{src},\xi^{l}_{tgt})=1/(1+e^{10\cdot(\bar{d}(\xi^{k}_{src},\xi^{l}_{tgt})-0.01)})$. Mini-batch construction then proceeds by (1) sampling a pair of trajectories $(\xi_{src},\xi_{tgt})$ with probability biased towards pairs exhibiting high similarity (i.e., low DTW distance) and (2) subsequently sampling individual transition tuples $(o_{src},x_{src},a_{src})$ and $(o_{tgt},x_{tgt},a_{tgt})$ from this selected, behaviorally similar trajectory pair.
Fine-grained temporal alignment, such as sampling around DTW-matched time steps, can optionally be employed here. We describe how the UOT loss (Eq. 3) is adapted with this new sampling procedure in the appendix.
@@ 4.4 Joint Co-Training Framework
Putting all components together, our final approach is a joint co-training framework where the visual feature encoder $f_{\phi}$ and the policy $\pi_{\theta}(a|z,x)$ are optimized concurrently. The Unbalanced Optimal Transport loss ($L_{\text{UOT}}$) serves as a regularization term, guiding $f_{\phi}$ to learn domain-invariant and action-relevant latent representations $z=f_{\phi}(o)$, while standard Behavior Cloning (BC) losses drive the policy learning. The overall training objective $L(f_{\phi},\pi_{\theta})=L_{\text{BC}}(f_{\phi},\pi_{\theta})+\lambda\cdot L_{\text{UOT}}(f_{\phi})$ combines these components, where $L_{\text{BC}}$ represents the combined behavior cloning losses calculated over both source ($D_{src}$) and target ($D_{tgt}$) datasets using a standard imitation loss (e.g., MSE). The hyper-parameter $\lambda>0$ balances feature alignment with policy imitation. The $L_{\text{UOT}}(f_{\phi})$ term is computed as defined in Equation 3, with mini-batches sampled with strategy described in Sec. 4.3. The overall training process is detailed in the appendix.
@@ 5.1 Experiment Setups
Environment setup. For the real-world experiments, we deploy our method on a Franka Emika Panda robotic arm, controlled at 20Hz using joint impedance control. Visual and point cloud observations are captured using an Intel RealSense D435 depth camera. Our simulation environments use the Robosuite Zhu et al. (2020) simulation framework. We calibrate the camera pose and camera intrinsics in simulation to match those obtained from the real-world setup to reduce the domain gap.
Simulation data. For simulation experiments, we begin by collecting 10 human demonstrations per task. Using MimicGen Mandlekar et al. (2023), we synthesize 200-1000 trajectories in the source domain, covering the full range of initial states (denoted as Source). In the target domain, we divide the reset region into two subregions: one is populated with 10 trajectories for training (denoted as Target), while the other remains completely held out from training (denoted as Target-OOD). This held-out subregion is used to evaluate each method’s generalization under Out-Of-Distribution (OOD) conditions.
@@
Observation modality and domain gaps. We evaluate our approach using two observation modalities: point clouds and RGB images. For point cloud observations, our method and the baselines adapt 3D Diffusion Policy (Ze et al., 2024) with a PointNet encoder (Qi et al., 2017). For RGB image observations, we use Diffusion Policy (Chi et al., 2023a) with a ResNet-18 encoder (He et al., 2016). To evaluate the generalization capabilities of different methods under visual domain shifts in simulation, we introduce several target domain variations: Viewpoint1-Point, Viewpoint3-Point, Perturbation-Point, Viewpoint-Image, and Texture-Image. Descriptions of each domain shift are provided in the appendix.
Baselines. We compare our method against the following baselines: MMD—minimizes the distance between the mean embeddings of source and target data (Tzeng et al., 2014); Co-training—trains the model using a mixed batch of source and target domain data, following the strategy proposed by (Wei et al., 2025; Maddukuri et al., 2025); Source-only—trains the model exclusively on data from the source domain, which in sim-to-real experiments corresponds to using only simulation data; Target-only—trains the model exclusively on data from the target domain, which in sim-to-real experiments trains with only real world data.
@@ 5.2 Cross-Domain Generalization Results
Our method effectively learns complex manipulation tasks in both simulation and the real world (H1). Experimental results show that our approach consistently matches or outperforms in terms of success rates all baselines across source and target domains in both simulated and real-world settings. On real-world tasks, our method achieves average success rates of 0.73 and 0.77 for image-based and point cloud-based policies, respectively. The Target-only baseline performs well in distribution but fails to generalize under domain shifts. The MMD baseline Tzeng et al. (2014) offers limited improvement by aligning global feature statistics, but its coarse alignment often disrupts task-relevant structure and harms source-domain performance. In contrast, by using Unbalanced Optimal Transport, our method performs selective, structure-aware alignment, avoiding spurious matches.
Our method generalizes to target domains only seen in simulation (H2). In Target-OOD scenarios, our method outperforms all baselines, underscoring the value of learning domain-invariant representations (see Tab. 2 and Tab. 4). While Co-training baselines Wei et al. (2025); Maddukuri et al. (2025) perform well when the target-domain training data overlaps with the evaluation region, they struggle to generalize when this overlap is absent. This limitation is especially evident under large domain shifts. For example, in the real-world BoxInBin and Stack tasks with novel textures or reset poses, our method achieves success rates of 0.7 and 0.4, respectively, using image-based observations—compared to just 0.3 and 0.0 for the Co-training baseline. These results highlight the shortcomings of relying purely on supervised target-domain data without explicitly addressing domain shift.
@@
Our method learns shared latent space that aligns simulation and real data. To better understand how the learned embeddings contribute to generalization, we visualize them using t-SNE Van der Maaten and Hinton (2008), as shown in Fig. 4(a). Blue and red correspond to features extracted from source and target domain observations, respectively. The left plot shows embeddings from the encoder trained with the Co-training baseline, while the right plot shows embeddings from our method. The visualization reveals that our approach leads to significantly better alignment between source and target distributions, highlighting its ability to learn domain-invariant representations that facilitate robust generalization. We also visualize the transport plan on a randomly sampled batch, along with the corresponding image observations from the source and target domains. As shown in Fig. 10 in the Appendix, the transport plan effectively aligns data points with similar states across domains.
@@ 6 Limitations and Conclusions
The main limitation of our work is that we only address sim-to-real visual observation gaps. Addressing action dynamics gaps remains future work. Because we use MimicGen Mandlekar et al. (2023) for automated simulation demonstration generation, we inherit its limitations, namely, our method primarily applies to tasks with prehensile and quasi-static interactions. We require a small number of real-world on-task demonstrations that are aligned with the simulated demonstrations. Future work involves relaxing this requirement, for example, by instead consuming unstructured real-world data, such as play data.
@@ 7 Acknowledgment
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The paper clearly states the claims made and contributions in the introduction and abstract.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitation of our method is discussed in the conclusion section.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification:
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The complete implementation detail is included in the Appendix
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The code and data are available on our project webpage.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We include full detail in Appendix
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: The main result of our paper requires substantial effort to execute each real world evaluation trial. Therefore, producing error bars is impractical. We provide detailed description on the evaluation conditions of our experiment.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide full detail on computation resource required for our experiments.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our experiment does not involve formal human study. Demonstration data is collected by authors with explicit consent.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: Our work is largely an algorithm for improving robot performance with simulation data. We do not expect immediate and direct societal impact of our work.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification:
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Our work is built on existing open source codebases, which we provide adequate citation and credit.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification:
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification:
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification:
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.